# Learning telic-controllable state representations

### v1

@@ 4 Discussion
We illustrated a novel approach to modelling purposeful behavior in bounded agents, based on the hypothesis that goals, defined as preferences over experience distributions, play a fundamental role in shaping state representations. Coupling together descriptive and normative aspects of learning models, our framing posits a granularity-complexity tradeoff as a theoretical grounding for modelling how human (and non-human) agents select which features of their environment to attend to (and at what resolution) and which to ignore (Niv et al., 2015; Langdon et al., 2019). While methods for goal-directed state abstraction have been previously proposed (Li et al., 2006; Abel et al., 2019; Shah et al., 2021; Kaelbling, 1993) our chief motivation is different, namely to provide a unified theoretical perspective on the role of goals in shaping state representation learning by complexity constrained cognitive agents. Our quantification of policy complexity follows previous work applying information theoretic principles in reinforcement learning (Rubin et al., 2012) and cognitive science (Amir et al., 2020). Notably, our complexity-granularity curves (Fig. 3) qualitatively resemble rate-distortion curves in information theory (Cover, 1999), suggesting a new interpretation of state representation learning via information theoretic lens (Arumugam & Van Roy, 2021). Finally, the duality between goals and state representations characterizing our approach may help address the thorny problem of goal formation: where do goals come from in the first place? Specifically, goals may be selected based on the properties of the state representations they produce; we accordingly hypothesize that bounded agents would prefer, all else being equal, goals that produce telic-controllable state representations, that balance environmental control with goal flexibility (cf. Klyubin et al. (2005)).
@@ Appendix B Pseudocode for telic-controllable state representation learning algorithm
Appendix C Dual-goal navigation task illustration
Figure 4: Dual goal navigation task: each tile shows 500 one-dimensional random walk trajectories of length $T=30$, generated by a Gaussian policy parameterized by the mean ($\mu$) and standard deviation ($\sigma$) of position update step (x-axis) across time (y-axis). Regions of interest $R$ and $L$ consist of line segments centered around $x_{R}=2$ and $x_{L}=-2$, shown as green and red lines respectively at $T=30$. Trajectories reaching one of the goals are plotted in the corresponding color, illustrating the relationship between policy parameters and goal reaching likelihoods. The default policy, $(\mu_{0},\sigma_{0})=(0,1)$, shown in the center gray tile, is equally likely to reach .
Appendix D Telic-complexity curves

### v2

@@ 4 Discussion
We illustrated a novel approach to modelling purposeful behavior in bounded agents, based on the hypothesis that goals, defined as preferences over experience distributions, play a fundamental role in shaping state representations. Coupling together descriptive and normative aspects of learning models, our framing posits a granularity-complexity tradeoff as a theoretical grounding for modelling how human, and non-human, agents select which features of their environment to attend to, and at what resolution, and which to ignore (Niv et al., 2015; Langdon et al., 2019). While several approaches for goal-directed state abstraction have been previously proposed (Li et al., 2006; Shah et al., 2021; Kaelbling, 1993; Zhang et al., 2020) our emphasis in this work is on principled theoretical perspective on the role of goals in shaping state representation learning by complexity constrained cognitive agents. Our quantification of policy complexity follows previous work applying information theoretic principles in reinforcement learning (Tishby & Polani, 2010; Rubin et al., 2012) and cognitive science (Amir et al., 2020; Lai & Gershman, 2024). Notably, our complexity-granularity curves (Fig. 4) qualitatively resemble rate-distortion curves in information theory (Cover, 1999), suggesting a new interpretation of state representation learning via information theoretic lens (Arumugam & Van Roy, 2021; Abel et al., 2019). Finally, the duality between goals and state representations characterizing our approach may help address the thorny problem of goal formation: where do goals come from in the first place? Specifically, goals may be selected based on the properties of the state representations they produce; we accordingly posit that bounded agents should prefer, all else being equal, goals producing telic-controllable state representations, that balance control over the environment with with behavioral adaptability (cf. Klyubin et al. (2005)).
@@ Appendix B Pseudocode for telic-controllable state representation learning algorithm
Appendix C Telic-complexity curves